Zeronknight (zeronknight):

ind the solution to the following inequality. x - 12 < 7 A. x < 19 B. x < 5 C. x > 5 D. x > 19

Zeronknight (zeronknight):

@Zarkon @pooja195

Zeronknight (zeronknight):

Hello?

pooja195 (pooja195):

simply add 12 to both sides

Zeronknight (zeronknight):

12+12=24 7+12=19

pooja195 (pooja195):

-12+12=0 2nd one is correct

pooja195 (pooja195):

So what would your answer be?

Zeronknight (zeronknight):

oh

Zeronknight (zeronknight):

A X < 19

pooja195 (pooja195):

Good

Zeronknight (zeronknight):

3000 medals in math..0-0

pooja195 (pooja195):